OpenStudy (anonymous):

2 square root 15 + square root 6 - 3 square root 45

OpenStudy (anonymous):

60 not 6

OpenStudy (pfenn1):

Is your expression this?\[2 \sqrt{15}+\sqrt{60} - 3 \sqrt{45}\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Hint:\[\Large 2 \sqrt{15}+\sqrt{60} - 3 \sqrt{45}=2\sqrt{3\cdot 5}+\sqrt{2^2 \cdot 3 \cdot 5}-3\sqrt{3^2\cdot 5}\]

OpenStudy (anonymous):

@yakeyglee how do you get the latex big like that ? :)

OpenStudy (anonymous):

I usually start with \Huge, \huge, \Large, or \large depending on how I'm feeling.

OpenStudy (anonymous):

haha cool, thx